# BackdoorMBTI: A Backdoor Learning Multimodal Benchmark Tool Kit for Backdoor Defense Evaluation

### v1

@@ Abstract.
In order to facilitate the research in multimodal backdoor, we introduce BackdoorMBTI, the first backdoor learning toolkit and benchmark designed for multimodal evaluation across three representative modalities from eleven commonly used datasets. BackdoorMBTI provides a systematic backdoor learning pipeline, encompassing data processing, data poisoning, backdoor training, and evaluation. The generated poison datasets and backdoor models enable detailed evaluation of backdoor defense methods. Given the diversity of modalities, BackdoorMBTI facilitates systematic evaluation across different data types. Furthermore, BackdoorMBTI offers a standardized approach to handling practical factors in backdoor learning, such as issues related to data quality and erroneous labels. We anticipate that BackdoorMBTI will expedite future research in backdoor defense methods within a multimodal context. Code is available at https://anonymous.4open.science/r/BackdoorMBTI-D6A1/README.md.
data poisoning, backdoor attack, backdoor defense, multimodal evaluation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Artificial intelligenceccs: Security and privacy
@@ 1. Introduction
As current efforts are primarily concentrated on designing efficient algorithms within specific domains such as computer vision, there is a pressing need for modality support. However, the existing benchmarks and defense methods are practically inadequate to meet such a need, primarily due to the diversity of modal data many real-world applications involve. This problem is caused by multiple factors. Firstly, the complexity of real-world data makes it challenging to conduct research and evaluate defense methods effectively in multimodal areas, resulting in uncertainties about their effectiveness. Secondly, the absence of a standardized evaluation baseline makes it hard to provide an objective evaluation of different algorithms. Lastly, current solutions often overlook empirical factors such as noise, undermining their overall efficacy in the real world. Nonetheless, addressing these issues is challenging. This is driven by the inherent complexity of the migration task, particularly within the framework of a multimodal application that necessitates support for diverse modalities and models. Additionally, attacks and defenses may involve numerous private parameters in their settings, making it difficult to evaluate them under a standard baseline.
@@
  • To enhance reproducibility and facilitate multimodal backdoor research, BackdoorMBTI offers an open-source backdoor learning framework that supports image, text, and audio tasks. We have developed a unified evaluation pipeline that is both user-friendly and easily extensible. Additionally, BackdoorMBTI provides the community with accessible poisoned datasets and models for defense evaluation. The source codes are available at https://anonymous.4open.science/r/BackdoorMBTI-D6A1/README.md.
@@ 2. Related Work
Table 2. An overview of 11 datasets included in BackdoorMBTI.
@@
  • *CelebA is a face attributes classification dataset, and all the face attributes are labeled in a bin format, in order to do the multi-class classification task on it, we follow the same settings in (Wang et al., 2022c).
@@ 2.1. Backdoor Attack
Existing backdoor attacks are typically categorized into three types, data poisoning attacks, training control attacks, and model modification attacks (Guo et al., 2022; Li et al., 2022). Data poisoning attacks involve the adversary manipulating the training data only (Gu et al., 2017; Chen et al., 2017; Nguyen and Tran, 2021; Liu et al., 2020; Li et al., 2021a; Liu et al., 2018b; Barni et al., 2019; Souri et al., 2022; Zeng et al., 2021b), while training control attacks allow the adversary to not only manipulate the data but also control the training process (Yao et al., 2019; Doan et al., 2021; Bagdasaryan and Shmatikov, 2021; Nguyen and Tran, 2020; Li et al., 2021f). Model-modified attacks, as described in (Bai et al., 2022; Tang et al., 2020; Qi et al., 2022), enable the adversary to manipulate the model directly.
@@ 2.2. Backdoor Defense
The existing defense methods can be categorized according to the machine learning lifecycle as follows:
@@
In our benchmark, we aim to encompass all the mentioned categories. Currently, at least one method has been implemented in each category, except for preprocessing and model-level detection, which will be supported in the future.
@@ 3. Supported Datasets
The BackdoorMBTI framework includes 11 datasets, as shown in Table 2, covering 8 different tasks. One significant reason for selecting these 11 datasets is their public availability, ensuring easy access. Additionally, these datasets are widely used in current backdoor learning research experiments. For each modality, we have implemented a commonly used model; for instance, ResNet for computer vision, BERT for natural language processing, and CNN for speech recognition. Other models can be extended easily in our benchmark.
@@ 4.1. Implemented Attack Algorithms
Table 3. The implemented attacks in BackdoorMBTI.
Modality Attack Visible Pattern Add SS*
@@
  • *sample specific, whether the trigger is sample specific.
@@ 4.2. Implemented Defense Algorithms
Table 4. The implemented defenses in BackdoorMBTI.
@@
  • a) Input: BM, Backdoor Model; CD, Clean Dataset; PD, Poison Dataset, a dataset including malicious backdoor data in it.
  • b) Stage: IT, In Training stage; PT, Post Training stage.
  • c) Output: CM, Clean Model, backdoor model after migration or a secure-training model; CD, Clean Dataset, a sanitized dataset; TP, Trigger Pattern, backdoor trigger pattern reversed by the defense.
In our multimodal framework, we have implemented 7 different backdoor defense methods. When selecting these defense methods, we first considered their theoretical applicability to multimodal scenarios. We also took into account the categories to which these defense methods belong and chose recent works as references.
@@ 5.1. Architecture Overview
4) Evaluation. The evaluation module takes a backdoor model and a curated test set as input and outputs performance metrics. The curated test set is specifically designed for evaluation purposes. It is generated using the data poisoning module with a poison ratio of 100%, wherein all instances with the attack target label are excluded. Detailed information about attack and defense performance metrics can be found in Section 6.1.
Table 5. The performance overview of backdoor attacks and defenses. CIFAR-10, SST-2, and SpeechCommands are used in the experiments for image, text, and audio modality separately.
@@ 5.2. Noise Generator Design
To address the authenticity challenge, we chose Gaussian noise as our primary noise generator due to its distribution and widespread use. For adaptability in text noise, we utilized an open-source noise generator called textnoisr  [^1] https://github.com/preligens-lab/textnoisr to introduce random modifications, deletions, and additions to the original text, simulating natural text noises. For controllability, we used noise ratios to adjust noise intensity in both image and audio data, and the character error rate to control noise levels in text.
@@ 6. Experiments
Figure 2. The accuracy comparison of various attack-defense pairs. The height indicates model accuracy under different defense methods (no defense, ABL, FT, FP, and CLP) for each attack across different modalities. Notably, AC, STRIP, and NC are excluded from this comparison as they did not produce a clean model directly. The asterisk denotes a value smaller than 0.001.
Figure 3. The ASR comparison of various attack-defense pairs. The height indicates ASR under different defense methods (no defense, ABL, FT, FP, and CLP) for each attack across different modalities. Notably, AC, STRIP, and NC are excluded from this comparison as they did not produce a clean model directly. The asterisk denotes a value smaller than 0.001.
@@ 6.1. Experiment Settings
Due to space limitations and training costs, four attacks (BadNets, BPP, SSBA, and WaNet) are selected in our experiments, more results can be accessed at https://anonymous.4open.science/r/BackdoorMBTI-D6A1/README.md. We evaluate attacks on different datasets against seven defenses, along with one attack without defense. Our default poisoning ratio is set at 10%.
@@
6.1.4. Metrics.
The metrics employed in evaluation are outlined as follows: Clean Accuracy (CAC) indicates the classification accuracy of the model when trained using a clean dataset. Backdoor Accuracy (BAC) reflects the classification accuracy when the model is trained using the backdoor dataset. Attack Success Rate (ASR) represents the rate of successful attacks, indicating when the model accepts a sample with a trigger and produces a targeted classification result. Robustness Accuracy (RAC) measures the robustness, demonstrating the ability of the model to provide correct classification results even with a trigger patch applied to the sample. For backdoor detection method, we used Detection Accuracy (DAC), recall (REC) and F1 score as metrics, DAC is the backdoor sample detection accuracy used for evaluating backdoor detection methods.
@@ 6.2. Overall Results (Q1)
All attacks after migration exhibit a significantly high attack success rate, exceeding 80% in general and surpassing 95% specifically for text. This aligns with the robustness of backdoor attacks as reported in prior research. Among the input filtering methods, AC (Chen et al., 2018) demonstrates promising results but maintains consistent effectiveness across all modalities at a low level. STRIP (Gao et al., 2019) achieves a near-perfect recall rate, often approaching 100%, and performs better on text compared to image and audio. ABL (Li et al., 2021b) performs exceptionally well on images, with a remarkable reduction in ASR, but falters in audio and text. As illustrated in Figure 4, FT (Liu et al., 2018a) demonstrates consistent performance across all modalities without compromising accuracy on the original task. FP (Liu et al., 2018a) and CLP (Zheng et al., 2022a), both based on pruning, encounter challenges due to their pruning operations are conducted on the batch normalization layer, which is absent in BERT, leading to their failure in text. However, FP outperforms CLP and exhibits effectiveness in audio. NC (Wang et al., 2019) excels in image defense but imposes input size constraints, limiting its applicability to text and audio without modification. Furthermore, ABL’s inability to identify backdoor data in UltraSonic (Koffas et al., 2022) attacks renders it ineffective against this particular threat. NC relies on reversing the trigger, but this reversed trigger fails to cause misclassification on normal data, explaining its failure against the BPP (Wang et al., 2022c) attack. However, it is imperative to emphasize the efficacy of NC in tackling noise conditions.
@@ 6.3. The Impact of Noise Factors (Q2)
One of our primary research questions involves investigating backdoor defense performance in a multimodal context within a real-world paradigm. We simulate real-world scenarios using Gaussian noise on data and mislabeling on labels. When employing these noise factors, we observe no significant decline in clean accuracy, with a decrease of around 3% falling within our expected range. As depicted in Figures 2 and 3, migrated attacks demonstrate a high success rate, indicating that noise factors do not adversely affect backdoor attacks. This is consistent with the robust nature of backdoor attacks, where the attachment of a backdoor trigger patch to an input sample leads to mispredictions regardless of the target label, even with noisy input.
While defense methods benefit from noise, multimodal defenses exhibit better migration results under noise. Our findings indicate a performance improvement in defense under noise, as shown by statistical analysis of experiment results. For mislabeled data, the improvement is 3.17%, and for noise data, it is 9.18%. The effectiveness of defense methods is contingent upon the quality of the backdoor model, training with noisy data yields a more robust model, making it easier to mitigate the impact of backdoor attacks. However, there is an exception where noise lowers defense performance in audio data. We believe this is explainable because audio data is concise and contains more compact information.

### v2

@@ Abstract.
In order to facilitate the research in multimodal backdoor, we introduce BackdoorMBTI, the first backdoor learning toolkit and benchmark designed for multimodal evaluation across three representative modalities from eleven commonly used datasets. BackdoorMBTI provides a systematic backdoor learning pipeline, encompassing data processing, data poisoning, backdoor training, and evaluation. The generated poison datasets and backdoor models enable detailed evaluation of backdoor defenses. Given the diversity of modalities, BackdoorMBTI facilitates systematic evaluation across different data types. Furthermore, BackdoorMBTI offers a standardized approach to handling practical factors in backdoor learning, such as issues related to data quality and erroneous labels. We anticipate that BackdoorMBTI will expedite future research in backdoor defense methods within a multimodal context. Code is available at https://github.com/SJTUHaiyangYu/BackdoorMBTI.
data poisoning; backdoor attack; backdoor defense; multimodal evaluation
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709385isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Artificial intelligenceccs: Security and privacy
@@ 1. Introduction
Modern backdoor attacks are not confined to images and have broadened their scope to include text and audio-related scenarios in the real world (Li et al., 2022). Some attacks even support multiple modalities simultaneously (Hammoud et al., 2023; Han et al., 2024). However, current efforts are primarily focused on designing efficient algorithms within specific domains such as computer vision. For example, highly integrated benchmarks like TrojanZoo are difficult to extend to new modalities. Although BackdoorBench supports both images and text, it separates the implementations of these two modalities due to differences in data processing, model loading, and training. These existing benchmarks are practically inadequate, highlighting an urgent need for modality support.
The aforementioned problem stems from multiple factors. Firstly, the complexity of real-world data makes it challenging to conduct research and evaluate defense methods effectively in multimodal areas, resulting in uncertainties about their effectiveness. Secondly, the absence of a standardized evaluation baseline makes it hard to provide an objective evaluation of different algorithms. Lastly, current solutions often overlook empirical factors such as noise, undermining their overall efficacy in the real world. Nonetheless, addressing these issues is challenging. This is driven by the inherent complexity of the migration task, particularly within the framework of a multimodal application that necessitates support for diverse modalities and models. Additionally, attacks and defenses may involve numerous private parameters in their settings, making it difficult to evaluate them under a standard baseline.
@@
  • To enhance reproducibility and facilitate multimodal backdoor research, BackdoorMBTI offers an open-source backdoor learning framework that supports image, text, and audio tasks. We have developed a unified evaluation pipeline that is both user-friendly and easily extensible. Additionally, BackdoorMBTI provides the community with accessible poisoned datasets and models for defense evaluation. The source codes, accompanied by user-friendly documents, are accessible at https://github.com/SJTUHaiyangYu/BackdoorMBTI.
@@ 2. Related Work
Table 2. An overview of 11 datasets included in BackdoorMBTI. *CelebA is a face attributes classification dataset, and all the face attributes are labeled in a bin format, in order to do the multi-class classification task on it, we follow the same settings in (Wang et al., 2022c).
@@ 2.1. Backdoor Attack
Existing backdoor attacks are typically categorized into three types, data poisoning attacks, training control attacks, and model modification attacks (Guo et al., 2022; Li et al., 2022). Data poisoning attacks involve the adversary manipulating the training data only (Gu et al., 2017; Chen et al., 2017; Nguyen and Tran, 2021; Liu et al., 2020; Li et al., 2021a; Liu et al., 2018b; Barni et al., 2019; Souri et al., 2022; Zeng et al., 2021b), while training control attacks allow the adversary to not only manipulate the data but also control the training process (Yao et al., 2019; Doan et al., 2021; Bagdasaryan and Shmatikov, 2021; Nguyen and Tran, 2020; Li et al., 2021f). Model modification attacks, as described in (Bai et al., 2022; Tang et al., 2020; Qi et al., 2022), enable the adversary to manipulate the model directly.
Backdoor attacks can also be categorized into various other types. For example, from the perspective of attack goals, adversaries aim to achieve higher stealthiness (Turner et al., 2019; Chen et al., 2017; Nguyen and Tran, 2021; Liu et al., 2020; Li et al., 2021a; Barni et al., 2019; Souri et al., 2022; Doan et al., 2021), pursue a higher attack success rate (Souri et al., 2022; Li et al., 2021a), or launch attacks without modifying labels (Turner et al., 2019; Liu et al., 2020; Li et al., 2021a; Barni et al., 2019; Souri et al., 2022). To enhance stealthiness, researchers are focusing on designing invisible patches that are imperceptible to human eyes when applied to image data. Regarding label modification, it can be divided into clean-label attacks, where the label remains unchanged, and dirty-label attacks. Additionally, there are other types of attacks, such as those with sample-specific dynamic patterns (Nguyen and Tran, 2021; Li et al., 2021a; Doan et al., 2021), where each data sample has a unique corresponding backdoor patch.
@@ 2.2. Backdoor Defense
Backdoor defense in multi-modal scenarios pertains to operations (e.g., detection, mitigation, or removal) targeting backdoor triggers across various modalities (e.g., images, text, audio). The existing defense methods can be categorized according to the machine learning lifecycle as follows:
@@
In our benchmark, we aim to encompass all the mentioned categories. Currently, at least one method has been implemented in each category, except for preprocessing, which will be supported in the future.
@@ 3. Supported Datasets and Models
The BackdoorMBTI framework includes 11 datasets, as shown in Table 2, covering 8 different tasks. One significant reason for selecting these 11 datasets is their public availability, ensuring easy access. Additionally, these datasets are widely used in current backdoor learning research experiments. For each modality, we have selected a commonly used model for experiments; for instance, ResNet for computer vision, BERT for natural language processing, and CNN for speech recognition. A total of 22 models have been implemented in our benchmark (refer to our repository page), with the flexibility to easily extend it with additional models.
@@ 4.1. Implemented Attack Algorithms
Table 3. The implemented attacks in BackdoorMBTI. Sample specific ($S$), whether the trigger is sample specific.
Modality Attack Visible Pattern Add $S$
@@ 4.2. Implemented Defense Algorithms
Table 4. The implemented defenses in BackdoorMBTI. a) Input: BM, Backdoor Model; CD, Clean Dataset; PD, Poison Dataset, a dataset including malicious backdoor data in it. b) Stage: IT, In Training stage; PT, Post Training stage. c) Output: CM, Clean Model, backdoor model after migration or a secure-training model; CD, Clean Dataset, a sanitized dataset; TP, Trigger Pattern, backdoor trigger pattern reversed by the defense.
@@
In our multimodal framework, we have implemented seven different backdoor defense methods. When selecting these defense methods, we first considered their theoretical applicability to multimodal scenarios from three perspectives: 1) general principle: for instance, STRIP uses entropy, which is applicable across data types. 2) common techniques, such as tuning and pruning, adapt across modalities. 3) practical observations, for example, CLP leverages minor perturbations (as backdoor models are sensitive to them) for detection. We also took into account the categories of these defense methods and chose recent works as references.
@@ 5.1. Architecture Overview
4) Evaluation. The evaluation module takes a backdoor model and a curated test set as input and outputs performance metrics. The curated test set is specifically designed for evaluation purposes. It is generated using the data poisoning module with a poison ratio of 100%, wherein all instances with the attack target label are excluded. Detailed information about attack and defense performance metrics can be found in Table 5.
Table 5. The performance overview of backdoor attacks and defenses. CIFAR-10, SST-2, and SpeechCommands are used in the experiments for image, text, and audio modality separately. Metrics used include Clean Accuracy (CAC) for clean model classification, Backdoor Accuracy (BAC) for backdoor model classification, Attack Success Rate (ASR) for successful triggered attacks, Robustness Accuracy (RAC) for correct classification despite trigger presence, and Detection Accuracy (DAC), recall (REC), and F1 score for evaluating backdoor detection methods.
@@ 5.2. Noise Generator Design
To address the authenticity challenge, we chose Gaussian noise as our primary noise generator due to its distribution and widespread use. For adaptability in text noise, we utilized a noise generator based on two open-source libraries (Félix Martel, 2023; Ma, 2019), which introduce character-level, word-level, and sentence-level perturbations to the original text, simulating natural text noises. Detailed information about the noise generator will be presented in Appendix A. For controllability, we employed noise ratios to adjust noise intensity in both image and audio data, and the character error rate to control noise levels in text.
@@ 6.1. Experiment Settings
Due to space limitations and training costs, four attacks (BadNets, BPP, SSBA, and WaNet) are selected in our experiments, more results can be accessed at https://github.com/SJTUHaiyangYu/BackdoorMBTI. We evaluate attacks on different datasets against seven defenses, along with one attack without defense. Our default poisoning ratio is set at 10%.
@@ 6.2. Overall Results (Q1)
Figure 2 provides a comparison of model accuracy under backdoor attacks and defenses, illustrating their usability visually. In the same context, Figure 3 displays the backdoor preservation metric ASR, facilitating the identification of effective defense methods. Moreover, Figure 4 illustrates the relationship between accuracy and ASR of backdoor defenses, with effective methods typically positioned in the top-left corner, indicating high accuracy and low ASR on sanitized models.
We employ accuracy, recall, and F1 scores as additional metrics for evaluating detection defense methods. Additionally, we have attempted to retrain the victim model using a sanitized dataset after detection. However, the retraining’s efficacy is quite low and lacks meaningful reference value. Model repair methods are evaluated using the same metrics as the attacks (BAC, ASR, RAC).
6.2.1. The performance of migrated attacks.
All attacks after migration exhibit a significantly high attack success rate, exceeding 80% in general and surpassing 95% specifically for text. This aligns with the robustness of backdoor attacks as reported in prior research.
6.2.2. The performance of migrated defenses.
As depicted in Table 5, AC, STRIP, FT, and FP consistently yield promising results across various experiments. One significant factor contributing to their success is that these defense methods are universal and widely applicable, thus demonstrating effectiveness in multimodal scenarios. Among the input filtering methods, AC (Chen et al., 2018) demonstrates promising results but maintains consistent effectiveness across all modalities at a low level. STRIP (Gao et al., 2019) achieves a near-perfect recall rate, often approaching 100%, and performs better on text compared to image and audio. ABL (Li et al., 2021b) performs exceptionally well on images, with a remarkable reduction in ASR, but falters in audio and text. As illustrated in Figure 4, FT (Liu et al., 2018a) demonstrates consistent performance across all modalities without compromising accuracy on the original task. FP (Liu et al., 2018a) and CLP (Zheng et al., 2022a), both based on pruning, encounter challenges due to their pruning operations are conducted on the batch normalization layer, which is absent in BERT, leading to their failure in text. However, FP outperforms CLP and exhibits effectiveness in audio. NC (Wang et al., 2019) excels in image defense but imposes input size constraints, limiting its applicability to text and audio without modification. Furthermore, ABL’s inability to identify backdoor data in UltraSonic (Koffas et al., 2022) attacks renders it ineffective against this particular threat. NC relies on reversing the trigger, but this reversed trigger fails to cause misclassification on normal data, explaining its failure against the BPP (Wang et al., 2022c) attack. However, it is imperative to emphasize the efficacy of NC in tackling noise conditions.
Figure 2. The accuracy comparison of various attack-defense pairs. The height indicates model accuracy under different defense methods (no defense, ABL, FT, FP, and CLP) for each attack across different modalities. Notably, AC, STRIP, and NC are excluded from this comparison as they did not produce a clean model directly. The asterisk denotes a value smaller than 0.001. Gaussian noise (mean 0, variance 1) and text noise (level 0.1) are used in the experiment.
Figure 3. The ASR comparison of various attack-defense pairs. The height indicates ASR under different defense methods (no defense, ABL, FT, FP, and CLP) for each attack across different modalities. Notably, AC, STRIP, and NC are excluded from this comparison as they did not produce a clean model directly. The asterisk denotes a value smaller than 0.001. Gaussian noise (mean 0, variance 1) and text noise (level 0.1) are used in the experiment.
6.2.3. The analysis of migrated defenses.
The findings in (Sha et al., 2022) demonstrate that fine-tuning effectively mitigates backdoors in models. This widely used technique is suitable for all deep learning scenarios, and given its model-agnostic and data-agnostic nature, it undoubtedly exhibits favorable behavior across various modalities.
STRIP (Gao et al., 2019), which relies on entropy theory and leverages the robustness characteristics of backdoors, serves as a fundamental and reasonable approach that performs well across modalities.
Owing to the strong correlation between the batch normalization layer and backdoor weights, FP (Liu et al., 2018a) and CLP (Zheng et al., 2022a) effectively mitigate backdoors. However, the capabilities of FP and CLP are limited by their dependency on the batch normalization layer. Future extensions to other layers and a clearer understanding of the relationship between model structure and backdoor will facilitate the further development of pruning-based methods. Nevertheless, it remains highly dependent on the model, specifically relying on the pruning decision algorithm.
The trigger reverse approach NC (Wang et al., 2019) holds promise for mitigating backdoor attacks. Its underlying assumption is universally applicable across modalities, and given its impressive performance in image data, it is reasonable to expect that NC will exhibit similar excellent performance in other modalities as well. As illustrated in (Sha et al., 2022), FT is highly effective in mitigating backdoors in models by adjusting their weights. As a widely used technique, it is applicable across various deep learning scenarios. Being both model-agnostic and data-agnostic, FT demonstrates reliable performance across different modalities.
@@ 6.3. The Impact of Noise Factors (Q2)
One of our primary research questions involves investigating backdoor defense performance in a multimodal context within a real-world paradigm. We simulate real-world scenarios using Gaussian noise on data and mislabeling on labels.
As depicted in Figures 2 and 3, migrated attacks demonstrate a high success rate, indicating that noise factors do not adversely affect backdoor attacks. This is consistent with the robust nature of backdoor attacks, where the attachment of a backdoor trigger patch to an input sample leads to mispredictions regardless of the target label, even with noisy input.
We conducted Mann-Whitney U tests to assess the detection accuracy (DAC) under three conditions: normal data, data noise, and label noise. For data noise, the p-value was 0.197, indicating no statistically significant difference compared to normal data. However, for label noise, the p-value was 0.012, demonstrating a significant improvement in DAC. Specifically, the average DAC was 62.8% under normal conditions, 65.5% in the presence of data noise, and 68.2% in the presence of label noise. These results suggest that defense methods benefit from noise, particularly label noise, and that multimodal defenses exhibit superior migration performance under such conditions.